The smallest angle in a triangle is half of the largest angle. The third angle is 10° less than the largest angle. Find the measure of all three angles.
Answer by Alan3354(69443)

The smallest angle in a triangle is half of the largest angle. The third angle is 10° less than the largest angle. Find the measure of all three angles.
------------
C = largest angle
A = C/2
B = C - 10
-----
A + B + C = 180
C/2 + (C-10) + C = 180
etc